1)Adenine is 30%.What would be percentage of guanine?
A)10% B)20% C)30% D)40%

2)In a DNA molecule cytosine is 18%.Percentage of adenine would be-
A)32% B)64% C)36% D)18%

3)Total amount of A and T in DNA is 45%.Amount of Guanine will be-
A)22.5% B)27.5% C)45% C)55%

4)In a DNA molecule Adenine is 31% and Guanine is 40%.What would be percentage of cytosine?
A)19% B)40% C)36% D)50%

Please explain easily.
1
Expert's answer
2014-08-12T05:12:17-0400
1) Adenine is 30%.What would be percentage of guanine?
A) 10% B)20% C)30% D)40%
Answer: B. The percentage of T isequal to the percentage of A, 30%. Total percentage of total C and G is
100-30-30=40, so the percentage of G is 40/2=20%.

2) In a DNA molecule cytosine is18%.Percentage of adenine would be-
A) 32% B)64% C)36% D)18%
Answer: A. The percentage of G isequal to the percentage of C, 18%. Total percentage of total A and T is
100-18-18=64, so the percentage of A is 64/2=32%.

3) Total amount of A and T in DNAis 45%.Amount of Guanine will be-
A)22.5% B)27.5% C)45% C)55%
Answer: B. Total amount of C and Gis 100-45=55%, the amount of G is 55/2=27.5%.

4) In a DNA molecule Adenine is 31% and Guanine is 40%.What would be percentage
of cytosine?
A)19% B)40% C)36% D)50%
Answer: B. The percentage of C is equal to the percentage of G, 40%.